# Exact and Linear Convergence for Federated Learning under Arbitrary Client Participation is Attainable

### v2

@@ E.3 Proof of the Convergence of SG-FOCUS ($\mu-$Strong Convexity)
$\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{r\tau+1}-{\boldsymbol{x}}_{r\tau}\|^{2}_{F}$
@@
E.4 Proof of the Convergence of SG-FOCUS (Non-Convexity with PL Assumption)
@@
E.5 Proof of the Convergence of SG-FOCUS (General Non-Convexity)
@@
$-\eta\left\langle\nabla F(\bar{x}_{r\tau+1}),\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau})\right\rangle$
@@
[…]
@@
$\frac{1}{R}\sum_{r=0}^{R-1}\mathbb{E}\hskip 1.42271pt\|\nabla F(\bar{x}_{r\tau+1})\|^{2}$
@@
$\hfill\qed$
Appendix F Supplementary Experiments for SG-FOCUS
F.1 Experiment Setup
To examine SG-FOCUS ’s performance under arbitrary client participation and the highly non-iid conditions, we compare SG-FOCUS with FedAU Wang and Ji (2023) and SCAFFOLD Karimireddy et al. (2020) on the image classification task by using the CIFAR10 Krizhevsky et al. (2009) dataset. The model we used is a three-layer convolutional neural network. As for our baselines, FedAU is a typical FL algorithm designed to tackle unknown client participation, and SCAFFOLD is a classic FL algorithm designed to deal with data heterogeneity. Therefore, it is reasonable to select these two FL algorithms as our baselines.
@@ F.1 Experiment Setup
F.2 Experiment Results of SG-FOCUS
From Figure 9, we observe that under highly heterogeneous data distribution ($\alpha=0.05$) and arbitrary client participation/sampling, our SG-FOCUS shows the best performance on convergence speed and test accuracy, compared to FedAU and SCAFFOLD. These results echo the view in our main paper: SG-FOCUS can jointly handle both data heterogeneity and arbitrary client participation.

### v3

@@ F.3 Proof of the Convergence of SG-FOCUS ($\mu$-Strong Convexity)
$\hskip-14.22636pt\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{r\tau+1}-{\boldsymbol{x}}_{r\tau}\|^{2}_{F}$
@@
F.4 Proof of the Convergence of SG-FOCUS (Non-Convexity with PL Assumption)
@@
F.5 Proof of the Convergence of SG-FOCUS (General Non-Convexity)
@@
$\hskip-14.22636pt-\eta\left\langle\nabla F(\bar{x}_{r\tau+1}),\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau})\right\rangle$
@@
$\;\;\;+\left(\frac{\eta}{2N}+\eta^{2}L\right)\mathbb{E}\hskip 1.42271pt\left\|\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}(\bar{x}_{r\tau+1})-\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau})\right\|^{2}$
@@
$\hskip-14.22636pt\frac{1}{R}\sum_{r=0}^{R-1}\mathbb{E}\hskip 1.42271pt\|\nabla F(\bar{x}_{r\tau+1})\|^{2}$
@@
[…]
Appendix G Supplementary Experiments for SG-FOCUS
G.1 Experiment Setup
To examine SG-FOCUS ’s performance under arbitrary client participation and the highly non-iid conditions, we compare SG-FOCUS with FedAU (Wang and Ji, 2024) and SCAFFOLD (Karimireddy et al., 2020) on the image classification task by using the CIFAR10 (Krizhevsky et al., 2009) dataset. The model we used is a three-layer convolutional neural network. As for our baselines, FedAU is a typical FL algorithm designed to tackle unknown client participation, and SCAFFOLD is a classic FL algorithm designed to deal with data heterogeneity. Therefore, it is reasonable to select these two FL algorithms as our baselines.
@@ G.1 Experiment Setup
G.2 Experiment Results of SG-FOCUS
From Figure 9, we observe that under highly heterogeneous data distribution ($\alpha=0.05$) and arbitrary client participation/sampling, our SG-FOCUS shows the best performance on convergence speed and test accuracy, compared to FedAU, MIFA, and SCAFFOLD. These results echo the view in our main paper: SG-FOCUS can jointly handle both data heterogeneity and arbitrary client participation.